# Modelling the Spread of New Information on Social Networks

### v3

@@ 6.2 Discussion
The gap between the two predictions is particularly pronounced for the model using message features M only, whose performance collapses for out-of-distribution prediction. This not only reveals the limited generalisation capability of models relying on message data, but also highlights the importance of conducting both in-distribution and out-of-distribution evaluations.
@@
The model using U-HM features, extracted from the text content of users’ historical messages, performs similarly to message-based model (M). Its out-of-distribution performance declines noticeably, though not as severely as M. This suggests that, beyond M, U-HM also contributes little to out-of-distribution prediction compared with features of user profiles (U-P) and user historical actions (U-HA).
@@
In contrast, three user-based models (U, U-P and U-HA) demonstrate significantly stronger predictive power, producing a performance as good as the model using ALL features for both in-distribution and out-of-distribution predictions. This is because these user-related features encode a user’s status and activity patterns that are independent from the latest trends on X and therefore can generalise across different topics.
6.2.2 The Decision Tree
Performance of the DT-ALL model, using all types of features, is higher that that of the NN-ALL model, using all types of data, in both in-distribution experiments (for mixed and individual hashtags) and out-of-distribution experiments.
@@ 6.2.2 The Decision Tree
This can perhaps be explained by the ability of decision trees to disregard uninformative features through their feature-splitting mechanism: at each node, the tree chooses the feature that best separates the classes, which not only reveals useful signals hidden in the data, but also effectively ignores weak or irrelevant features.
6.2.3 It is who we are, not what we see.
Our results show that reposting behaviour for previously unseen topics can be predicted by users’ profile and past behaviour, and is largely content-agnostic. In other words, who the users are plays a more central role than what information they are exposed to.
This aligns with recent LLM-agent simulations [55, 56] showing that echo chambers and extreme voices emerge irrespective of recommendation strategies. This observation at the population level echos our results on reposting prediction at the level of individual users. That is, if exchange of new information between individual users are largely content-agnostic, then the overall social dynamics in a population would be insensitive, or even immune, to recommendation algorithms.
@@
Figure 2: In-distribution and out-of-distribution predictions by Decision Tree (DT) models using different types of features. User-related features U includes U-P, U-HA and U-HM.
@@ 7 Conclusion
Reposting prediction is a fundamental problem in modelling information diffusion. Existing works largely focused on textual data and only considered in-distribution evaluation. Our experimental results show that while message-based methods perform well in in-distribution scenarios, they show poor performance in out-of-distribution tasks when evaluated on repostings associated with previously unseen topics. We demonstrate that models incorporating user-related data always outperform message-based models. In particular, for out-of-distribution prediction, adding user-related data can significantly increase F1 score from 0.117 to 0.705.
Our findings suggest that a significant component of users’ reposting decisions, particularly for previously unseen topics, can be predicted based on users’ profile and past behaviour, and is largely content-agnostic.
We also argue that to ensure comprehensive evaluation while addressing practical needs, prediction works on social media should always include both in-distribution and out-of-distribution assessments.
@@ Acknowledgments
Appendix A Data Ethics
In our research, we collected data including messages and user profiles from publicly available sources (X platform, formerly Twitter). All data utilized in our study has been strictly limited to academic research purposes. The data has not been used to target or identify any individual personally. We have taken rigorous measures to ensure the anonymity and privacy of the data subjects. We have not disclosed any of the data to third parties.
Appendix B Hyper-parameters of Decision Tree Models
We searched for the optimal hyper-parameters of the DT models for each experiment setting. Searched values are provided in Table 6, where values marked with * are most frequently used. The parameter scale_pos_weight was only considered for imbalanced datasets (1:5 and 1:10).
Table 6: Hyper-parameters of Decision Tree models
@@ Appendix B Hyper-parameters of Decision Tree Models
[1.00*, 3.25, 5.50, 7.75, 10.00] (1:10 dataset)
Table 7: In-Distribution Prediction for Individual Hashtags using Temporal Split
Model DT-ALL DT-U DT-M
@@
Model NN-ALL NN-U NN-M BERT
Dataset F1 Score (mean $\pm$ std)
[…]
Appendix C Additional Temporal Prediction
@@
Table 7 reports the additional temporal prediction results, i.e. in-distribution prediction results for individual hashtags using the temporal split. For each hashtag, we divide the dataset into 13 temporal windows with equal numbers of positive instances. The first 3 windows are used for training and the next for testing, rolling forward 1 window at each time step. The latest 10% of the training data is used for validation. Ten sequential predictions will be conducted. The results and conclusions are consistent with those presented in the main paper.
Appendix D Full Lists of Features
@@
The following tables show the full lists of Message (M) features, User Profile (U-P) features, and User Historical Action (U-HA) features. The full list of User Historical Message (U-HM) features is in Table 1.
@@
The following language models and tools are used for calculating message features.
@@
  • Topic features: Twitter-roBERTa-base  [^4] https://huggingface.co/cardiffnlp, and Latent Dirichlet Allocation model from scikit-learn.
@@
  • Language features: language-tool-python  [^5] https://github.com/jxmorris12/language_tool_python, TextBlob  [^6] https://github.com/sloria/textblob, and a Logistic Regression model trained on a public data source  [^7] https://www.kaggle.com/datasets/crowdflower/twitter-user-gender-classification.
@@
  • Readability features: a Python implementation of several readability measures  [^8] https://github.com/andreasvc/readability/.
@@
  • Sentiment features: VADER module from NLTK  [^9] https://www.nltk.org/index.html.
@@
  • Emotion & Hate-Speech features: A Transformer-based library for SocialNLP tasks  [^10] https://github.com/pysentimiento/pysentimiento.
Table 8: Message (M) Features
@@
Topic (39) M_TopicM1 Likelihood that the message belongs to a specific topic in the 19-topic scheme 1.
@@
M_TopicG1 Likelihood that the message belongs to a specific topic in the 6-topic scheme 2.
@@
M_TopicLDA1 Likelihood that the message belongs to 1 of the 10 topics 3 identified by the LDA model.
@@
Language (10) M_CharNum The number of characters of the message.
M_WordNum The number of words of the message.
@@
M_Grammar2 Scores for grammatical and spelling correctness of the whole message.
M_Polarity The orientation of the expressed tone of the message.
M_Subjectivity A measure of personal opinion and factual information contained in the message.
M_Irony A measure of how ironic the message is.
M_Offensive A measure of how offensive the message is.
@@
Readability (11) M_Readability1 Different measures of how difficult the message in English is to understand.
@@
M_Readability10 Count of words in the message that consists of three or more syllables.
M_Readability11 Count of words in the message that are not on the list used in DaleChallIndex.
Sentiment (5) M_Sentiment1 Negative sentiment score of the message.
M_Sentiment2 Neutral sentiment score of the message.
M_Sentiment3 Positive sentiment score of the message.
@@
Emotion (8) M_Emotion1 The probability that the message expresses a specific type of emotion.
@@
M_EmotionMain The most likely emotion expressed by the message.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of the message.
M_Hate2 A measure of hatefulness of the message.
M_Hate3 A measure of whether the message is targeting an individual (or entity).
M_HsNum The # of the above types of hate speech (above given thresholds) detected in the message.
Hashtag (1) M_hashtag The categorical label of the hashtag under study contained in the message.
1 The 19 topic-scheme includes: arts&culture, business&entrepreneurs, celebrity&pop_culture, diaries&daily_life, family, fashion&style, film_tv&video, fitness&health, food&dining, gaming, learning&educational, music, news&social_concern, other_hobbies, relationships, science&technology, sports, travel&adventure, youth&student_life. 2 The 6 topic-scheme includes: arts&culture, business&entrepreneurs, pop_culture, daily_life, sports&gaming, science&technology. 3 The top 3 keywords for each topic discovered by LDA model: 1) blm, amp, last; 2) nhs, covid, go; 3) john, mp, dido; 4) nhs, ios, one; 5) new, china, america; 6) loveisland, davide, think; 7) energy, brexit, energyprices; 8) avengers, marvel, news; 9) monkeypox, ukraine, russian; 10) covid, youtube, supercup.
Table 9: User Profile (U-P) Features
@@
U-P_R_TweetNum The number of total messages that R has posted since registration.
@@
U-P_R_SpreadActivity The ratio of the # of R’s messages to the maximal # of messages among all users.
@@
U-P_R_TweetNumDay The number of total messages normalized by the age of this account.
@@
Table 10: User Historical Action (U-HA) Features
@@
U-HA_R_TweetNum The number of R’s historical messages.
U-HA_R_TweetPercent Percentage of original-tweets/retweets/quotes/replies in R’s historical messages.
@@
U-HA_R_InteractivePer Percentage of reposts in R’s historical messages.
U-HA_R_AverageInterval The average time interval (days) between the posting of R’s historical messages.
U-HA_R_RetweetedRate Average number of retweets/quotes/replies/likes received by R’s historical messages.
@@
U-HA_RS_Mention The number of times that R mentioned S in R’s historical messages.
U-HA_RS_MentionPer Percentage of messages in which R mentioned S in R’s historical messages.
U-HA_SR_Mention The number of times that S mentioned R in S’s historical messages.
U-HA_SR_MentionPer Percentage of messages in which S mentioned R in S’s historical messages.
Interaction U-HA_RS_RepostLatency Latency (days) from the creation time of the message to the reposting.
@@
U-HA_SR_PathWidth The cosine distance between the message’s topic vector and the Topic-Oriented

### v4

@@ 6 Discussion
The gap between the two predictions is particularly pronounced for the model using post features M only, whose performance collapses for out-of-distribution prediction. This not only reveals the limited generalisation capability of models relying on post data, but also highlights the importance of conducting both in-distribution and out-of-distribution evaluations.
@@
The model using U-HM features, extracted from the text content of users’ historical posts, performs similarly to the post-based model (M). Its out-of-distribution performance declines noticeably, though not as severely as M. This suggests that, beyond M, U-HM also contributes little to out-of-distribution prediction compared with features of user profiles (U-P) and user historical actions (U-HA).
@@
In contrast, three user-based models (U, U-P, and U-HA) demonstrate significantly stronger predictive power, producing a performance as good as the model using ALL features for both in-distribution and out-of-distribution predictions. This is because these user-related features encode a user’s status and activity patterns that are independent from the latest trends on X and therefore can generalise well across different topics.
6.2 DT models vs. NN models
@@
Performance of the DT-ALL model, using all types of features, is higher than that of the NN-ALL model, using all types of data, in both in-distribution experiments (for mixed and individual hashtags) and out-of-distribution experiments. This can perhaps be explained by the ability of decision trees to disregard uninformative features through their feature-splitting mechanism: at each node, the tree chooses the feature that best separates the classes, which not only reveals useful signals hidden in the data, but also effectively ignores weak or irrelevant features.
6.3 Users Shape Reposting Behaviour
Our results suggest that a significant component of reposting behaviour for previously unseen topics can be predicted from user profiles and past behaviours, and is largely content-agnostic. This supports the observation in [10] that cognitive and behavioural mechanisms are not peripheral details but fundamental determinants of information diffusion.
This, however, does not mean that reposting behaviour is independent of post content. Rather, our results suggest that both user-related and post-related factors contribute to reposting behaviour, as evidenced by the strong performance of the in-distribution predictions based on both user-related and post-related data.
The key difference is that content-related factors are inherently difficult to generalise to previously unseen topics due to semantic shifts; whereas user-related factors are largely content-agnostic, and therefore can maintain their predictive power when moving from in-distribution to out-of-distribution predictions.
6.4 Future Work
A recent work based on LLM-agent simulations [31, 43] showed that users can still form polarized echo chambers on social networks even without content-based recommendation algorithms. Future work should study how user characteristics, post content, and algorithmic exposure jointly influence information diffusion.
As multimodal content becomes increasingly common on social media, future work should study pictures and videos embedded in posts.
@@ 6.4 Future Work
Figure 2: In-distribution and out-of-distribution predictions by Decision Tree (DT) models using different types of features, where U = U-P + U-HA + U-HM.
@@ 7 Conclusion
Reposting prediction is a fundamental problem in modelling information diffusion. Existing works largely focused on textual data and only considered in-distribution evaluation. Our experimental results show that while post-based methods perform well in in-distribution scenarios, they show poor performance in out-of-distribution tasks when evaluated on repostings associated with previously unseen topics. We demonstrate that models incorporating user-related data always outperform post-based models. In particular, for out-of-distribution prediction, adding user-related data can significantly increase the F1 score from 0.117 to 0.705.
Our findings suggest that a significant component of users’ reposting decisions, particularly for previously unseen topics, can be predicted based on user profiles and past behaviours, and is largely content-agnostic.
We also argue that to ensure comprehensive evaluation while addressing practical needs, prediction works on social networks should always include both in-distribution and out-of-distribution assessments.
@@ Acknowledgments
Appendix A Ethics Statement
In our research, we collected data including posts and user profiles from publicly available sources (X platform, formerly Twitter) under the platform’s terms of service. All data utilized in our study has been strictly limited to academic research purposes. The data has not been used to target or identify any individual personally. We have taken rigorous measures to ensure the anonymity and privacy of the data subjects. We have not disclosed any of the data to third parties.
Appendix B Computational Resources
All experiments in this study were conducted on a Linux-based workstation, with an AMD Ryzen 9 9950X CPU, 64 GB of system memory, and an NVIDIA GeForce RTX 4090 GPU. The operating system is Rocky Linux 9.6.
Our models are computationally efficient and compatible with standard GPU-equipped servers. No conclusions in this study depend on specialized hardware.
Appendix C Hyper-parameters of Decision Tree Models
We searched for the optimal hyper-parameters of the Decision Tree (DT) models for each experiment setting. The hyper-parameters were determined on the validation set. Searched values are provided in Table 6, where values marked with * are most frequently used. The parameter scale_pos_weight was only considered for imbalanced datasets (1:5 and 1:10).
@@ Appendix C Hyper-parameters of Decision Tree Models
[1.00*, 3.25, 5.50, 7.75, 10.00]
(1:10 dataset)
Table 6: Hyper-parameters of Decision Tree models
Model DT-ALL DT-U DT-M NN-ALL NN-U NN-M BERT
@@
Table 7: In-distribution prediction for individual hashtags using temporal split.
Appendix D Calculation of Mean and Standard Deviation
For each hashtag, we compute the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ of the F1 scores across the 10 cross-validation folds. We then report the overall mean $\bar{\mu}$ and standard deviation $\bar{\sigma}$ across all 14 hashtags by treating the results from the 14 hashtags as equally weighted components of a mixture distribution.
$\bar{\mu}=\frac{1}{14}\sum_{i=1}^{14}\mu_{i}$ (2)
Appendix E Additional Results on a General 1:5 Dataset
In addition to the main experiments reported in the paper, where negative instances are constructed by selecting the five most similar non-reposted cases for each positive instance, we also conduct supplementary experiments on a more general 1:5 dataset, which serve as a robustness check and evaluate our models under a more general sampling strategy.
For convenience, we first define the set of posts in our raw data as $\mathcal{P}=\{(M_{h},U_{S,t_{s}})\}$ and the set of reposts $\mathcal{RP}=\{(M_{h},U_{S,t_{s}},U_{R,t_{r}})\}$ [^4] Note: Reposts whose original posts are absent from the raw data are excluded.. We further denote $\mathcal{U}=\{U_{t_{u}}\}$ as the set of users and their associated activity timestamps, which includes all $U_{S,t_{s}}$ from $\mathcal{P}$ and $U_{R,t_{r}}$ from $\mathcal{RP}$.
In the supplementary experiments, a negative instance $(M^{*}_{h},U^{*}_{S,t_{s}},U^{*}_{t_{u}})$ is generated by randomly pairing a post $(M^{*}_{h},U^{*}_{S,t_{s}})\in\mathcal{P}$ and a user $U^{*}_{t_{u}}\in\mathcal{U}$, on the condition that $(M^{*}_{h},U^{*}_{S,t_{s}},U^{*}_{t_{u}})\notin\mathcal{RP}$ and $t_{u}>t_{s}$. These conditions reflect the most fundamental causal requirements for a negative instance: a user $U^{*}_{t_{u}}$ can only repost a post $M^{*}_{h}$ if $U^{*}_{t_{u}}$ is active after $M^{*}_{h}$ is created, and no reposting event between $U^{*}_{t_{u}}$ and $M^{*}_{h}$ is observed in the data. Therefore, a negative instance is generated independently of a specific positive instance and without similarity constraints, which is more general than the strategy used in the main experiments.
The prediction results are shown in Table 8. The results and conclusions are consistent with those presented in the main paper.
Model DT-ALL DT-U DT-M
Experiment F1 Score
In-distribution (mixed) 0.916 0.907 0.775
In-distribution (individual) 0.901 0.894 0.740
Out-of-distribution 0.802 0.792 0.101
Table 8: Prediction results on the general 1:5 dataset. We focus on the Decision Tree (DT) models given their superior performance in the main experiments. Evaluations are conducted under three settings: (1) in-distribution (mixed) using data from all hashtags; (2) in-distribution (individual) using data from each individual hashtag, reported as the average across all individual hashtags; and (3) out-of-distribution, where the model is trained on 13 hashtags and tested on 1 remaining hashtags.
Appendix F Additional Results for Temporal Prediction
@@
Table 7 reports the additional temporal prediction results, i.e., in-distribution prediction results for individual hashtags using the temporal split. For each hashtag, we divide the dataset into 13 temporal windows with equal numbers of positive instances. The first 3 windows are used for training and the next for testing, rolling forward 1 window at each time step. The latest 10% of the training data is used for validation. Ten sequential predictions will be conducted. The results and conclusions are consistent with those presented in the main paper.
Appendix G Full Lists of Features
@@
The following tables show the full lists of Post (M) features, User Profile (U-P) features, and User Historical Action (U-HA) features. The full list of User Historical Post (U-HM) features is in Table 1.
@@
The following language models and tools are used for calculating post features.
@@
  • Topic features: Twitter-roBERTa-base 555 https://huggingface.co/cardiffnlp, and Latent Dirichlet Allocation model from scikit-learn.
@@
  • Language features: language-tool-python 666 https://github.com/jxmorris12/language_tool_python, TextBlob 777 https://github.com/sloria/textblob, and a Logistic Regression model trained on a public data source 888 https://www.kaggle.com/datasets/crowdflower/twitter-user-gender-classification.
@@
  • Readability features: a Python implementation of several readability measures 999 https://github.com/andreasvc/readability/.
@@
  • Sentiment features: VADER module from NLTK 101010 https://www.nltk.org/index.html.
@@
  • Emotion & Hate-Speech features: A Transformer-based library for SocialNLP tasks 111111 https://github.com/pysentimiento/pysentimiento.
@@
Topic (39) M_TopicM1 Likelihood that the post belongs to a specific topic in the 19-topic scheme 1.
@@
M_TopicG1 Likelihood that the post belongs to a specific topic in the 6-topic scheme 2.
@@
M_TopicLDA1 Likelihood that the post belongs to 1 of the 10 topics 3 identified by the LDA model.
@@
Language (10) M_CharNum The number of characters of the post.
M_WordNum The number of words of the post.
@@
M_Grammar2 Scores for grammatical and spelling correctness of the whole post.
M_Polarity The orientation of the expressed tone of the post.
M_Subjectivity A measure of personal opinion and factual information contained in the post.
M_Irony A measure of how ironic the post is.
M_Offensive A measure of how offensive the post is.
@@
Readability (11) M_Readability1 Different measures of how difficult the post in English is to understand.
@@
M_Readability10 Count of words in the post that consists of three or more syllables.
M_Readability11 Count of words in the post that are not on the list used in DaleChallIndex.
Sentiment (5) M_Sentiment1 Negative sentiment score of the post.
M_Sentiment2 Neutral sentiment score of the post.
M_Sentiment3 Positive sentiment score of the post.
@@
Emotion (8) M_Emotion1 The probability that the post expresses a specific type of emotion.
@@
M_EmotionMain The most likely emotion expressed by the post.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of the post.
M_Hate2 A measure of hatefulness of the post.
M_Hate3 A measure of whether the post is targeting an individual (or entity).
M_HsNum The # of the above types of hate speech (above given thresholds) detected in the post.
Hashtag (1) M_hashtag The categorical label of the hashtag under study contained in the post.
Table 9: Post (M) Features. 1 The 19 topic-scheme includes: arts&culture, business&entrepreneurs, celebrity&pop_culture, diaries&daily_life, family, fashion&style, film_tv&video, fitness&health, food&dining, gaming, learning&educational, music, news&social_concern, other_hobbies, relationships, science&technology, sports, travel&adventure, youth&student_life. 2 The 6 topic-scheme includes: arts&culture, business&entrepreneurs, pop_culture, daily_life, sports&gaming, science&technology. 3 The top 3 keywords for each topic discovered by LDA model: 1) blm, amp, last; 2) nhs, covid, go; 3) john, mp, dido; 4) nhs, ios, one; 5) new, china, america; 6) loveisland, davide, think; 7) energy, brexit, energyprices; 8) avengers, marvel, news; 9) monkeypox, ukraine, russian; 10) covid, youtube, supercup.
@@
U-P_R_TweetNum The number of total posts that R has posted since registration.
@@
U-P_R_SpreadActivity The ratio of the # of R’s posts to the maximal # of posts among all users.
@@
U-P_R_TweetNumDay The number of total posts normalized by the age of this account.
@@
Table 10: User Profile (U-P) Features
@@
U-HA_R_TweetNum The number of R’s historical posts.
U-HA_R_TweetPercent Percentage of original-tweets/retweets/quotes/replies in R’s historical posts.
@@
U-HA_R_InteractivePer Percentage of reposts in R’s historical posts.
U-HA_R_AverageInterval The average time interval (days) between the posting of R’s historical posts.
U-HA_R_RetweetedRate Average number of retweets/quotes/replies/likes received by R’s historical posts.
@@
U-HA_RS_Mention The number of times that R mentioned S in R’s historical posts.
U-HA_RS_MentionPer Percentage of posts in which R mentioned S in R’s historical posts.
U-HA_SR_Mention The number of times that S mentioned R in S’s historical posts.
U-HA_SR_MentionPer Percentage of posts in which S mentioned R in S’s historical posts.
Interaction U-HA_RS_RepostLatency Latency (days) from the creation time of the post to the reposting.
@@
U-HA_SR_PathWidth The cosine distance between the post’s topic vector and the Topic-Oriented